The geometrical length of a bar magnet having half magnetic length 5 centimetre is

Grade:11

2 Answers

Arun
5 years ago
The magnetic length of a bar magnet is nearly 0.84 times that of the geometric length(Physical length).
 
In question, magnetic length will be = 2*5 = 10 cm.
Hence 
Magnetic length = 0.84 * geometric length
Geometric length = 10/0.84 = 11.905 cm
Khimraj
4 years ago
Geometric length: Geometric length is the measure of total length of magnet. A bar magnet consist of two equal and opposite magnetic pole separated by a small distance. Poles are not exactly at the ends. The shortest distance between two poles is called effective length and is less than its geometric length.
Magnetic length = 0.84 * geometric length
Geometric length = 2*5/0.84 = 11.905 cm